*** Please use TI-83+ calculator to solve this problem ***

You wish to test the following claim (HaHa) at a significance level of α=0.02α=0.02.

      Ho:p1=p2Ho:p1=p2
      Ha:p1<p2Ha:p1<p2

You obtain 465 successes in a sample of size n1=592n1=592 from the first population. You obtain 593 successes in a sample of size n2=687n2=687 from the second population. For this test, you should NOT use the continuity correction, and you should use the normal distribution as an approximation for the binomial distribution.

What is the test statistic for this sample? (Report answer accurate to three decimal places.)
test statistic = Incorrect

What is the p-value for this sample? (Report answer accurate to four decimal places.)
p-value = Incorrect

The p-value is...

  • less than (or equal to) αα
  • greater than αα

Correct



This test statistic leads to a decision to...

  • reject the null
  • accept the null
  • fail to reject the null

Correct



As such, the final conclusion is that...

  • There is sufficient evidence to warrant rejection of the claim that the first population proportion is less than the second population proportion.
  • There is not sufficient evidence to warrant rejection of the claim that the first population proportion is less than the second population proportion.
  • The sample data support the claim that the first population proportion is less than the second population proportion.
  • There is not sufficient sample evidence to support the claim that the first population proportion is less than the second population proportion.
Answer #1

(a)

n1 = 592

p1 = 465/592 = 0.7855

n2 = 687

p2 = 593/687 =0.8632

nln2

595 + 591 = 0.8272 592 +687

Q = 1 - P = 0.1728

SE -VPQ(1/nl1/n2)

V/0.8272 × 0.1728(1/592 1/687) 0.02 12

Test statistic is:

Z = (0.7855 - 0.8632)/0.0212 = - 3.665

So,

test statistic is:

- 3.665

(b)

To find P - Value:

In TI-83+ calculator:

Select 5:1-PropZTest

Enter above statistical data.

Select Calculate.

We get:

p = 0.0001

(c)

Correct option:

less than or equal to alpha

(d)

Correct option:

reject the null

(e)

correct option:

The sample datasupport the claim that the first population proportion is less than the second population proportion.